OpenStudy (anonymous):

Using your function, explain to the Martians how to solve for f(3). Show your work and explain each step using complete sentences. my function F(x)= 4x + 8

OpenStudy (anonymous):

thanks for your help

OpenStudy (anonymous):

please help

OpenStudy (imstuck):

When you have a function, f(x), that means that the function is dependent upon whatever that x is. So when you have a function, f(x), and you are told that x = 3, or f(3), you determine the value of the function when you replace x with 3. That works like this:\[f(x)=4x+8\]\[f(3)=4(3)+8\]\[f(3)=12+8\]\[f(3)=20\]

OpenStudy (imstuck):

Tell your Martian friend "good luck"! lol

OpenStudy (anonymous):

thx